Consider a machine that produces metal pieces which are cylindrical in shape. We select
a sample of nine pieces and measure the diameters:
1.01; 0.97; 1.03; 1.04; 0.99; 0.98; 0.99; 1.01; 1.03
The sample and sample standard deviation are x̄ = 1.00556 and s = 0.02455, respectively.
Give a 95% confidence interval for the true mean diameter, assume that the
population is normal.
Answer #1

Solution :

Given that,

\bar x =1.00556

s = 0.02455

n = 9

Degrees of freedom = df = n - 1 = 9 - 1 = 8

At 95% confidence level the t is ,

\alpha  = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,8 =2.306

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.306 * (0.02455 / \sqrt 9)

= 0.01887

Margin of error = 0.01887

The 99% confidence interval estimate of the population mean is,

\bar x - E <\mu  < \bar x + E

1.00556 - 0.01887 < \mu <1 .00556 + 0.01887

0.9867 < \mu < 1.0244

(2.70, 3.70 )